OpenStudy (vshiroky):

Graph the equation. Be sure to simplify the equation into standard form Ax+By=C before graphing it. 3x+4y-3=-15

OpenStudy (vshiroky):

This one I am unsure of how to do..

jimthompson5910 (jim_thompson5910):

First add 3 to both sides 3x+4y-3=-15 3x+4y-3+3=-15+3 3x+4y = -12 Now the equation is in Ax+By = C form

jimthompson5910 (jim_thompson5910):

One way to graph this is to find the x and y intercepts To find the x-intercept, plug in y = 0 and solve for x 3x+4y = -12 3x+4(0) = -12 3x+0 = -12 3x = -12 x = -12/3 x = -4 So the x-intercept is the point (-4,0)

jimthompson5910 (jim_thompson5910):

The y-intercept is found the same way, but plugging in x = 0 and solving for y this time 3x+4y = -12 3(0)+4y = -12 0+4y = -12 4y = -12 y = -12/4 y = -3 telling us that the point (0,-3) is also on this line and this point is the y-intercept

OpenStudy (vshiroky):

Ok... Would the second one be (0,-3)?

jimthompson5910 (jim_thompson5910):

now plot the two points (-4,0) and (0,-3) and draw a straight line through them

jimthompson5910 (jim_thompson5910):

that is correct

OpenStudy (vshiroky):



jimthompson5910 (jim_thompson5910):

That looks good to me

OpenStudy (vshiroky):

Yay!! Thank you yet again haha.

jimthompson5910 (jim_thompson5910):

sure thing, glad to be of help